From the records of the Public Health Division, the percentages of ?A, ?B, ?O, and ??AB blood types are, respectively, 40%, 20%, 30% and 10%.

If two individuals are picked at random from this population, assuming independence, find the probability that

(a) they are both of blood type ?A.

(b) they are both of the same blood type.

(c) neither is of blood type ?A.

(d) one is of blood type ?A and the other is of blood type ?O.

(e) they are both of different blood types.
Answer #1

If A and B are independent then

P(A B) =P(AB) = P(A) *P(B)

A) P(AA)=0.4*0.4=0.16

B) Prob (both blood type same) =0.30

C) Prob (neither of blood type A) =0.6*0.6=0.36

D) Prob (one of blood type A and other of blood type O) =0.24

E) Prob (both of different blood type) =0.70

Solution file is attached go through it

Thanks